# Behavior Synthesis via Contact-Aware Fisher Information Maximization

### v1

@@ III-A Parameter Learning via Maximum a-posteriori Estimate
[…]
@@ III-B Fisher Information
Note that the Fisher Interestingly, the Fisher information matrix also has a unique interpretation as a lower-bound on the uncertainty of the posterior in maximum likelihood problem.
@@ IV Contact-Aware Experimental Design
s.t. $\begin{cases}x_{0}\text{ given }&\text{ (initial cond.) }\\ x_{t+1}=f_{\theta}(x_{t},u_{t},\lambda_{t})&\text{ (dynamics model) }\\ \lambda_{t}\in\mathcal{C}_{\theta}(x_{t})&\text{ (contact constraints) }\\ \theta\in\Theta&\text{ (feasible parameter) }\end{cases}$
@@ V-A Contact-Seeking Behavior Synthesis
[…]
@@ V-B Closing the Loop
Figure 6: Contact Excitations Improves Parameter Learning. We compare our experimental design approach with a baseline similar to works in [25] for all robot scenarios. The baseline implements a belief-space planning controller that reasons about contact naïvely. We evaluate each robot scenario over a finite number of experiments. Here, we model the parameter error as %$\%\mathrm{err}=100(\hat{\theta}-\theta^{*})/\theta^{*}$, where $\hat{\theta}$ is the current estimate and $\theta^{*}$ is the true parameter. By maximizing the contact-aware Fisher information, we see that the robot seeks contacts that are information-rich about the unknown parameters of interest, and thus converging closer to the true parameter values $\theta^{*}$, with resulting higher Fisher information gains.
@@ 4. Q4)
In Fig.7, we demonstrate the robustness of our approach as a function of parameter initialization. We observe that our approach is highly robust to initial priors within $\sim$30for hefting, and even more robust for other scenarios. We also expand on the remarks in [8] that estimating multiple parameters that are co-dependent to one another makes the estimation unidentifiable, unless we make assumptions of the contact interactions (e.g. bounding contact impulses), provided within the problem definition in Eq. 8 in the paper.
@@ Appendix B Implementation Details
$\begin{split}\left(||\mathcal{F}||^{2}||\mathcal{F}^{-1}||^{2}-\left(n-2\right)\right)^{2}\geq\left(\kappa\left(\mathcal{F}\right)+\kappa^{-1}\left(\mathcal{F}\right)\right)^{2}&\\ \geq\left\{\begin{array}[]{c}n\ \mathrm{even},\frac{4}{n^{2}}||\mathcal{F}||^{2}||\mathcal{F}^{-1}||^{2}\\ n\ \mathrm{odd},\frac{4}{n^{2}-1}\left(||\mathcal{F}||^{2}||\mathcal{F}^{-1}||^{2}-1\right)\end{array}\right.\end{split}$ (17)
where $\mathcal{F}\in\mathbb{R}^{n\times n}$, and the Schur matrix norm $||\mathcal{F}||=\mathrm{tr}\left(\mathcal{F}^{T}\mathcal{F}\right)$, for a nonsingular matrix $\mathcal{F}$. By maximizing over the trace of the Fisher information matrix in Eq. 7, we thus maximize the bounds of the condition number.
@@
Friction Estimation Experiments. The Franka end effector is free to move within 2 dimensional domain $\mathcal{W}=[0,1.0]\times[0.5,1.0]\ \mathrm{m}$ with respect to a local wall frame coordinate system at the base of the wall, with bounded joint velocity constraints $v_{ee}\in[-1.0,1.0]\ \mathrm{m/s}$ in both x and y directions. The robot starts at initial end effector position $p_{ee,0}=[0.5,0.8]\ \mathrm{m}$ in wall frame coordinates (oriented $0.8\ \mathrm{m}$ in the y direction from the root of the Franka, with its x axis pointing away from the wall) with zero initial velocity. The robot estimates the coefficient of friction between its end effector and the wall, with true parameter $\mu=0.4$. We initialize the variance of the mass parameter as $\Sigma_{\theta}=1$. The contact model parameters in this scenario is shown in Table I.
[…]

### v2

@@ III-A Parameter Learning via Maximum a-posteriori Estimate
$p(\theta|\mathcal{D})=\frac{p(\mathcal{D}|\theta)p(\theta)}{\int_{\Theta}p(\mathcal{D}|v)p(v)dv}.$ (1)
@@ III-B Fisher Information
Interestingly, the Fisher information matrix also has a unique interpretation as a lower-bound on the uncertainty of the posterior in maximum likelihood problem.
@@ IV Contact-Aware Experimental Design
[…]
@@ V-A Contact-Seeking Behavior Synthesis
$\text{ s.t. }\begin{cases}x_{0}\text{ (given)}\\ x_{t+1}=f_{\theta}(x_{t},u_{t},\lambda_{t}),\\ \lambda_{t}\in\mathcal{C}_{\theta}(x_{t})\\ y_{t}=g_{\theta}(x_{t})\end{cases}$
@@ V-B Closing the Loop
Figure 6: Contact Excitations Improves Parameter Learning. We compare our experimental design approach with a baseline similar to works in [25] for all robot scenarios. The baseline implements a belief-space planning controller that reasons about contact naïvely. We evaluate each robot scenario over a finite number of experiments. Here, we model the parameter error as $\%\mathrm{err}=100(\hat{\theta}-\theta^{*})/\theta^{*}$, where $\hat{\theta}$ is the current estimate and $\theta^{*}$ is the true parameter. By maximizing the contact-aware Fisher information, we see that the robot seeks contacts that are information-rich about the unknown parameters of interest, and thus converging closer to the true parameter values $\theta^{*}$, with resulting higher Fisher information gains.
@@ 4. Q4)
In Fig.7, we demonstrate the robustness of our approach as a function of parameter initialization. We observe that our approach is highly robust to initial priors within $\sim$30% for hefting, and even more robust for other scenarios. We also expand on the remarks in [8] that estimating multiple parameters that are co-dependent to one another makes the estimation unidentifiable, unless we make assumptions of the contact interactions (e.g. bounding contact impulses), provided within the problem definition in Eq. 8 in the paper.
@@ Appendix B Implementation Details
[…]
where $\mathcal{F}\in\mathbb{R}^{n\times n}$, and the Schur matrix norm $\mathcal{F}||=\mathrm{tr}\left(\mathcal{F}^{T}\mathcal{F}\right)$, for a nonsingular matrix $\mathcal{F}$. By maximizing over the trace of the Fisher information matrix in Eq. 7, we thus maximize the bounds of the condition number.
@@
Friction Estimation Experiments. The Franka end effector is free to move within 2 dimensional domain $\mathcal{W}=[0,1.0]\times[0.5,1.0]\ \mathrm{m}$ with respect to a local wall frame coordinate system at the base of the wall, with bounded joint velocity constraints $v_{ee}\in[-1.0,1.0]\ \mathrm{m/s}$ in both x and y directions. The robot starts at initial end effector position $p_{ee,0}=[0.5,0.8]\ \mathrm{m}$ in wall frame coordinates (oriented 0.8m0.8\ $\mathrm${m} in the y direction from the root of the Franka, with its x axis pointing away from the wall) with zero initial velocity. The robot estimates the coefficient of friction between its end effector and the wall, with true parameter $\mu=0.4$. We initialize the variance of the mass parameter as $\Sigma_{\theta}=1$. The contact model parameters in this scenario is shown in Table I.
Material Property Estimation Experiments. A robot gripper enters a pregrasped state, where the gripper surface is positioned $0.4$ cm away from the surface of the ball object with stiffness coefficient $K=800\ \textrm{N/m}$ N/m, and dampening coefficient $C=10\ \textrm{Ns/m}$. The gripper is free to move within gripper velocity constraints and prismatic joint constraints as per an Aloha 2 system. We initialize the variance of the mass parameter as $\Sigma_{\theta}=[100.0\ (\mathrm{N/m})^{2},10.0\ (\mathrm{Ns/m})^{2}]$.